# MaskGCT: Zero-Shot Text-to-Speech with Masked Generative Codec Transformer

### v1

@@ Abstract
Nowadays, large-scale text-to-speech (TTS) systems are primarily divided into two types: autoregressive and non-autoregressive. The autoregressive systems have certain deficiencies in robustness and cannot control speech duration. In contrast, non-autoregressive systems require explicit prediction of phone-level duration, which may compromise their naturalness. We introduce the Masked Generative Codec Transformer (MaskGCT), a fully non-autoregressive model for TTS that does not require precise alignment information between text and speech. MaskGCT is a two-stage model: in the first stage, the model uses text to predict semantic tokens extracted from a speech self-supervised learning (SSL) model, and in the second stage, the model predicts acoustic tokens conditioned on these semantic tokens. MaskGCT follows the mask-and-predict learning paradigm. During training, MaskGCT learns to predict masked semantic or acoustic tokens based on given conditions and prompts. During inference, the model generates tokens of a specified length in a parallel manner. We scale MaskGCT to a large-scale multilingual dataset with 100K hours of in-the-wild speech. Our experiments demonstrate that MaskGCT achieves superior or competitive performance compared to state-of-the-art zero-shot TTS systems in terms of quality, similarity, and intelligibility while offering higher generation efficiency than diffusion-based or autoregressive TTS models. Audio samples are available at https://maskgct.github.io/.
@@ 1 Introduction
In recent years, large-scale zero-shot text-to-speech (TTS) systems (Kharitonov et al., 2023; Wang et al., 2023; Łajszczak et al., 2024; Kim et al., 2024; Peng et al., 2024; Anastassiou et al., 2024) have achieved significant improvements by scaling data and model size, including both autoregressive (AR) (Kharitonov et al., 2023; Wang et al., 2023; Łajszczak et al., 2024; Kim et al., 2024; Peng et al., 2024; Anastassiou et al., 2024; Shen et al., 2023; Ju et al., 2024; Le et al., 2024; Jiang et al., 2023) and non-autoregressive (NAR) models (Shen et al., 2023; Ju et al., 2024; Le et al., 2024; Jiang et al., 2023). However, these systems still exhibit some problems. Currently, AR-based TTS systems typically quantize speech into discrete tokens and then use language models to sequentially generate these tokens, which offer diverse prosody but also suffer from problems such as poor robustness and slow inference speed. NAR-based models, typically based on diffusion (Ju et al., 2024; Shen et al., 2023), flow matching (Le et al., 2024), or GAN (Jiang et al., 2023), require text and speech alignment information as well as the prediction of phone-level duration, resulting in a complex pipeline and producing more standardized but less natural speech.
@@
Recently, masked generative transformers, a class of generative models, have achieved significant results in the fields of image (Chang et al., 2022, 2023; Li et al., 2023a), video Yu et al. (2023a, b), and audio (Garcia et al., 2023; Li et al., 2024; Ziv et al., 2024) generation, demonstrating potential comparable to or superior to autoregressive models or diffusion models. These models employ a mask-and-predict training paradigm and utilize iterative parallel decoding during inference. Some previous works have attempted to introduce masked generative models into the field of TTS. SoundStorm (Borsos et al., 2023) was the first attempt to use a masked generative transformer to predict multi-layer acoustic tokens extracted from SoundStream, conditioned on speech semantic tokens; however, it still needs to receive the semantic tokens of an AR model as input. Thus, SoundStorm is more of an acoustic model that converts semantic tokens into acoustic tokens and does not fully utilize the powerful generative potential of masked generative models. NaturalSpeech 3 (Ju et al., 2024) decomposes speech into discrete token sequences representing different attributes through special designs and generates tokens representing different attributes through masked generative models. However, NaturalSpeech 3 still requires speech-text alignment information to predict the phone-level duration.
@@
In this work, we propose MaskGCT, a fully non-autoregressive model for text-to-speech synthesis that uses masked generative transformers without requiring text-speech alignment information. MaskGCT is a two-stage system; both stages are trained using the mask-and-predict learning paradigm. The first stage, the text-to-semantic (T2S) model, predicts masked speech semantic representation tokens within an in-context learning paradigm, using text token sequences and prompt speech semantic token sequences as the prefix, without requiring precise alignment information between text and speech. The second stage, the semantic-to-acoustic (S2A) model, utilizes semantic tokens to predict masked acoustic tokens extracted from an RVQ-based speech codec with prompt acoustic tokens. During inference, MaskGCT can generate speech of various specified lengths with high quality and naturalness, requiring only 25 iteration steps for using text to predict semantic tokens, which is significantly fewer than AR-based models. Another key aspect of our method is that we train a VQVAE (Van Den Oord et al., 2017) to quantify self-supervised speech learning features, rather than using k-means to obtain semantic tokens as in previous work. This approach minimizes the information loss of semantic features even with a single codebook. We also explore the scalability of our methods in tasks such as speech translation, voice conversion, emotion control, and speech content editing, demonstrating the potential of MaskGCT as a foundational model for speech generation.
@@
Our experiments demonstrate that MaskGCT has achieved performance comparable to or superior to that of existing models in terms of speech quality, similarity, prosody, and intelligibility. Specifically, (1) It achieves comparable or better quality and naturalness than the ground truth speech across three benchmarks (LibriSpeech, SeedTTS test-en, and SeedTTS test-zh) in terms of CMOS. (2) It achieves human-level similarity between the generated speech and the prompt speech, with improvements of +0.017, -0.002, and +0.027 in SIM-O and +0.156, +0.319 and +0.244 in SMOS for LibriSpeech, SeedTTS test-en, and SeedTTS test-zh, respectively. (3) It achieves comparable intelligibility in terms of word error rate (WER) across the three benchmarks and demonstrates stability within a reasonable range of speech duration, which also indicates the diversity and controllability of the generated speech.
@@
The contribution of our proposed MaskGCT can be summarized as follows: (1) We propose an innovative system for zero-shot TTS, which is based on non-autoregressive masked generative transformers. We also introduce a speech discrete semantic representation based on training a VQVAE on speech self-supervised representations. (2) MaskGCT achieves human-level similarity, naturalness, and intelligibility by scaling data to 100K hours of in-the-wild speech. MaskGCT also demonstrates the ability to control the total duration of generated speech, showcasing the controllability and diversity of speech produced by MaskGCT. (3) We investigate the scalability of MaskGCT across various tasks, including speech translation, voice conversion, emotion control, and speech content editing, utilizing zero-shot learning or post-training methods. This exploration underscores the potential of MaskGCT as a foundational model for speech generation.
@@ 2 Related Work
Large-scale TTS. Traditional TTS systems (Ren et al., 2020, 2019; Tan et al., 2024; Wang et al., 2017; Kim et al., 2021) are trained to generate speech from a single speaker or multiple speakers using hours of high-quality transcribed training data. Modern large-scale TTS systems (Kharitonov et al., 2023; Wang et al., 2023; Łajszczak et al., 2024; Kim et al., 2024; Peng et al., 2024; Anastassiou et al., 2024) aim to achieve zero-shot TTS (synthesizing speech for unseen speakers with speech prompts) by scaling both the model and data size. These systems can be mainly divided into AR-based and NAR-based categories. For AR-based systems: SpearTTS (Kharitonov et al., 2023) utilizes three AR models to predict semantic tokens from text, coarse-grained acoustic tokens from semantic tokens, and fine-grained acoustic tokens from coarse-grained tokens. VALLE (Wang et al., 2023) predicts the first layer of acoustic tokens extracted from EnCodec (Défossez et al., 2022) using an AR LLM, and the final layers with a NAR LLM. VoiceCraft (Peng et al., 2024) employs a single AR model to predict multi-layer acoustic tokens in a delayed pattern (Copet et al., 2024). BASETTS (Łajszczak et al., 2024) predicts novel speech codes extracted from WavLM features and uses a GAN model for waveform reconstruction. For NAR-based systems: NaturalSpeech 2 (Shen et al., 2023) employs latent diffusion to predict the latent representations from a codec model (Zeghidour et al., 2021). VoiceBox (Le et al., 2024) uses flow matching and in-context learning to predict mel-spectrograms, following a similar framework. MegaTTS (Jiang et al., 2023) utilizes a GAN to predict mel-spectrograms, while an AR model predicts phone-level prosody codes. NaturalSpeech 3 (Ju et al., 2024) employs a unified framework based on discrete diffusion models to predict discrete representations of different speech attributes. However, these NAR systems need to predict phoneme-level duration, leading to a complex pipeline and more standardized generative results. SimpleSpeech (Yang et al., 2024), DiTTo-TTS (Lee et al., 2024), and E2 TTS (Eskimez et al., 2024) are also NAR-based models that do not require precise alignment information between text and speech, nor do they predict phoneme-level duration. We discuss these concurrent works in Appendix L.
@@
Masked Generative Model. Masked generative transformers, a class of generative models, achieve significant results and demonstrate potential comparable to or superior to that of autoregressive models or diffusion models in the fields of image (Chang et al., 2022, 2023; Lezama et al., 2022; Li et al., 2023a), video (Yu et al., 2023a, b), and audio (Borsos et al., 2023; Garcia et al., 2023; Li et al., 2024; Ziv et al., 2024) generation. MaskGIT (Chang et al., 2022) is the first work to use masked generative models for both unconditional and conditional image generation. Subsequently, Muse (Chang et al., 2023) leverages rich text to achieve high-quality and diverse text-to-image generation within the same framework. MAGVIT-v2 (Yu et al., 2023b) employs masked generative models with the novel lookup-free quantization (LFQ), outperforming diffusion models in image and video generation. Recently, some efforts have been made to adapt masked generative models to the field of audio generation. SoundStorm (Borsos et al., 2023) takes in the semantic tokens from AudioLM and utilizes this generative paradigm to generate tokens for a neural audio codec (Zeghidour et al., 2021). VampNet (Garcia et al., 2023) and MAGNeT (Ziv et al., 2024) apply masked generative models for music and audio generation, while MaskSR (Li et al., 2024) extends these models for speech restoration.
@@
Discrete Speech Representation. Speech representation is a crucial aspect of speech generation. Early works (Ren et al., 2019; Wang et al., 2017) typically utilized mel-spectrograms as the modeling target. Recently, some large-scale TTS systems (Wang et al., 2023; Ju et al., 2024) have shifted to using discrete speech representations. Discrete speech representation can be primarily divided into two types: semantic discrete representation and acoustic discrete representation. Semantic discrete representations are mainly extracted from various speech SSL models (Chung et al., 2021; Hsu et al., 2021; Chen et al., 2022) using quantization methods such as k-means. Acoustic discrete representations, on the other hand, are usually obtained by training a VQGAN model (Van Den Oord et al., 2017) with the goal of waveform reconstruction, as seen in speech neural codecs (Défossez et al., 2022; Zeghidour et al., 2021; Kumar et al., 2024; Yang et al., 2023). Semantic discrete representation typically shows a stronger correlation with text, whereas acoustic discrete representation more effectively reconstructs audio. Consequently, some two-stage TTS models predict both semantic and acoustic tokens. FACodec (Ju et al., 2024) is a novel speech codec that disentangles speech waveforms into subspaces of different attributes, including content, prosody, timbre, and acoustic details.
@@ 3.2 Model Overview
Following (Betker, 2023; Borsos et al., 2023; Wang et al., 2023), MaskGCT is a two-stage TTS system. The first stage uses text to predict speech semantic representation tokens, which contain most information of content and partial information of prosody. The second stage model is trained to learn more acoustic information. Unlike previous works (Betker, 2023; Wang et al., 2023; Borsos et al., 2023; Kharitonov et al., 2023) use an autoregressive model for the first stage, MaskGCT utilizes the non-autoregressive masked generative modeling paradigm for both the two stages without any alignment information between text and speech: (1) For the first stage model, we trained a model to learn $p_{\theta_{\text{s1}}}(\mathbf{S}|\mathbf{S}_{t},(\mathbf{S}^{p},\mathbf{P}))$, where $\mathbf{S}$ is the speech semantic representation token sequence obtained from a speech semantic representation codec (we introduce in 3.3), $\mathbf{S}^{p}$ is the prompt semantic token sequence, and $\mathbf{P}$ is the text token sequence. $\mathbf{S}^{p}$ and $\mathbf{P}$ are the condition for the first stage model. (2) The second stage model is trained to learn $p_{\theta_{\text{s2}}}(\mathbf{A}|\mathbf{A}_{t},(\mathbf{A}^{p},\mathbf{S}))$, where $\mathbf{A}$ is the multi-layer acoustic token sequence from a speech acoustic codec like (Zeghidour et al., 2021; Défossez et al., 2022). Our second stage model is similar to SoundStorm (Borsos et al., 2023). The overview of MaskGCT is presented in Figure 1. We give more details about the four parts in the following sections.
@@
Figure 1: An overview of our MaskGCT system. MaskGCT consists of four main parts: (1) a speech semantic representation codec converts speech to semantic tokens; (2) a text-to-semantic model predicts semantic tokens with text and prompt semantic tokens; (3) a semantic-to-acoustic model predicts acoustic tokens conditioned on semantic tokens; (4) a speech acoustic codec reconstructs speech waveform from acoustic tokens.
3.3 Speech Semantic Representation Codec

### v2

@@ Abstract
The recent large-scale text-to-speech (TTS) systems are usually grouped as autoregressive and non-autoregressive systems. The autoregressive systems implicitly model duration but exhibit certain deficiencies in robustness and lack of duration controllability. Non-autoregressive systems require explicit alignment information between text and speech during training and predict durations for linguistic units (e.g. phone), which may compromise their naturalness. In this paper, we introduce Masked Generative Codec Transformer (MaskGCT), a fully non-autoregressive TTS model that eliminates the need for explicit alignment information between text and speech supervision, as well as phone-level duration prediction. MaskGCT is a two-stage model: in the first stage, the model uses text to predict semantic tokens extracted from a speech self-supervised learning (SSL) model, and in the second stage, the model predicts acoustic tokens conditioned on these semantic tokens. MaskGCT follows the mask-and-predict learning paradigm. During training, MaskGCT learns to predict masked semantic or acoustic tokens based on given conditions and prompts. During inference, the model generates tokens of a specified length in a parallel manner. Experiments with 100K hours of in-the-wild speech demonstrate that MaskGCT outperforms the current state-of-the-art zero-shot TTS systems in terms of quality, similarity, and intelligibility. Audio samples are available at https://maskgct.github.io/.
@@ 1 Introduction
In recent years, large-scale zero-shot text-to-speech (TTS) systems Kharitonov et al. (2023); Wang et al. (2023); Łajszczak et al. (2024); Kim et al. (2024); Peng et al. (2024); Anastassiou et al. (2024); Shen et al. (2023); Ju et al. (2024); Le et al. (2024); Jiang et al. (2023) have achieved significant improvements by scaling data and model sizes, including both autoregressive (AR) Kharitonov et al. (2023); Wang et al. (2023); Łajszczak et al. (2024); Kim et al. (2024); Peng et al. (2024); Anastassiou et al. (2024) and non-autoregressive (NAR) models Shen et al. (2023); Ju et al. (2024); Le et al. (2024); Jiang et al. (2023). However, both AR-based and NAR-based systems still exhibit some shortcomings. In particular, AR-based TTS systems typically quantize speech into discrete tokens and then use decoder-only models to autoregressively generate these tokens, which offer diverse prosody but also suffer from problems such as poor robustness and slow inference speed. NAR-based models, typically based on diffusion Ju et al. (2024); Shen et al. (2023), flow matching Le et al. (2024), or GAN Jiang et al. (2023), require explicit text and speech alignment information as well as the prediction of phone-level duration, resulting in a complex pipeline and producing more standardized but less diverse speech.
@@
Recently, masked generative transformers, a class of generative models, have achieved significant results in the fields of image Chang et al. (2022, 2023); Li et al. (2023a), video Yu et al. (2023a, b), and audio Garcia et al. (2023); Li et al. (2024); Ziv et al. (2024) generation, demonstrating potential comparable to or superior to autoregressive models or diffusion models. These models employ a mask-and-predict training paradigm and utilize iterative parallel decoding during inference. Some previous works have attempted to introduce masked generative models into the field of TTS. SoundStorm Borsos et al. (2023a) was the first attempt to use a masked generative transformer to predict multi-layer acoustic tokens extracted from SoundStream, conditioned on speech semantic tokens; however, it needs to receive the semantic tokens of an AR model as input. Thus, SoundStorm is more of an acoustic model that converts semantic tokens into acoustic tokens and does not fully utilize the powerful generative potential of masked generative models. NaturalSpeech 3 Ju et al. (2024) decomposes speech into discrete token sequences representing different attributes through special designs and generates tokens representing different attributes through masked generative models. However, it still needs speech-text alignment supervision and phone-level duration prediction.
@@
In this work, we propose MaskGCT, a fully non-autoregressive model for text-to-speech synthesis that uses masked generative transformers without requiring text-speech alignment supervision and phone-level duration prediction. MaskGCT is a two-stage system, both stages are trained using the mask-and-predict learning paradigm. The first stage, the text-to-semantic (T2S) model, predicts masked semantic tokens with in-context learning, using text token sequences and prompt speech semantic token sequences as the prefix, without explicit duration prediction. The second stage, the semantic-to-acoustic (S2A) model, utilizes semantic tokens to predict masked acoustic tokens extracted from an RVQ-based speech codec with prompt acoustic tokens. During inference, MaskGCT can generate semantic tokens of various specified lengths with a few iteration steps given a sequence of text. In addition, we train a VQ-VAE Van Den Oord et al. (2017) to quantize speech self-supervised learning embedding, rather than using k-means to extract semantic tokens that is common in previous work. This approach minimizes the information loss of semantic features even with a single codebook. We also explore the scalability of our methods beyond the zero-shot TTS task, such as speech translation (cross-lingual dubbing), speech content editing, voice conversion, and emotion control, demonstrating the potential of MaskGCT as a foundational model for speech generation. Table 1 shows a comparison between MaskGCT and some previous works.
Table 1: A comparison between MaskGCT and existing systems. “Model” stands for modeling method and “Rep.” stands for the representation used. MaskGCT uses masked generative modeling for acoustic and semantic tokens (“A.” stands for acoustic, “S.” stands for semantic, “F.” stands for factorized tokens used in NaturalSpeech 3). MaskGCT implicitly models duration (“Imp. Dur.”) and allows flexible control over the total length of generated speech (“Len. Ctrl”). MaskGCT supports various speech generation tasks.
System Model Rep. Imp. Dur. Len. Ctrl. ZS TTS CL TTS Dubbing Edit
VALL-E Autoregressive A. Tokens
NaturalSpeech 2 Diffusion A. Features
VoiceBox Diffusion A. Features
VoiceCraft Autoregressive A. Tokens
NaturalSpeech 3 Masked Generative F. Tokens
MaskGCT Masked Generative S.&A. Tokens
@@
Our experiments demonstrate that MaskGCT has achieved performance comparable to or superior to that of existing models in terms of speech quality, similarity, prosody, and intelligibility. Specifically, (1) It achieves comparable or better quality and naturalness than the ground truth speech across three benchmarks (LibriSpeech, SeedTTS test-en, and SeedTTS test-zh) in terms of CMOS. (2) It achieves human-level similarity between the generated speech and the prompt speech, with improvements of +0.017, -0.002, and +0.027 in SIM-O and +0.28, +0.32 and +0.25 in SMOS for LibriSpeech, SeedTTS test-en, and SeedTTS test-zh, respectively. (3) It achieves comparable intelligibility in terms of WER across the three benchmarks and demonstrates stability within a reasonable range of speech duration, which also indicates the diversity and controllability of the generated speech.
@@
In summary, we propose a non-autoregressive zero-shot TTS system based on masked generative transformers and introduce a speech discrete semantic representation by training a VQ-VAE on speech self-supervised representations. Our system achieves human-level similarity, naturalness, and intelligibility by scaling data to 100K hours of in-the-wild speech, while also demonstrating high flexibility, diversity, and controllability. We investigate the scalability of our system across various tasks, including cross-lingual dubbing, voice conversion, emotion control, and speech content editing, utilizing zero-shot learning or post-training methods. This showcases the potential of our system as a foundational model for speech generation.
@@ 2 Related Work
Large-scale TTS. Traditional TTS systems Ren et al. (2020, 2019); Tan et al. (2024); Wang et al. (2017); Kim et al. (2021) are trained to generate speech from a single speaker or multiple speakers using hours of high-quality transcribed training data. Modern large-scale TTS systems Kharitonov et al. (2023); Wang et al. (2023); Łajszczak et al. (2024); Kim et al. (2024); Peng et al. (2024); Anastassiou et al. (2024) aim to achieve zero-shot TTS (synthesizing speech for unseen speakers with speech prompts) by scaling both the model and data size. These systems can be mainly divided into AR-based and NAR-based categories. For AR-based systems: SpearTTS Kharitonov et al. (2023) utilizes three AR models to predict semantic tokens from text, coarse-grained acoustic tokens from semantic tokens, and fine-grained acoustic tokens from coarse-grained tokens. VALL-E Wang et al. (2023) predicts the first layer of acoustic tokens extracted from EnCodec Défossez et al. (2022) using an AR codec language model, and the final layers with a NAR model. VoiceCraft Peng et al. (2024) employs a single AR model to predict multi-layer acoustic tokens in a delayed pattern Copet et al. (2024). BASETTS Łajszczak et al. (2024) predicts novel speech codes extracted from WavLM features and uses a GAN model for waveform reconstruction. For NAR-based systems: NaturalSpeech 2 Shen et al. (2023) employs latent diffusion to predict the latent representations from a codec model Zeghidour et al. (2021). VoiceBox Le et al. (2024) uses flow matching and in-context learning to predict mel-spectrograms. MegaTTS Jiang et al. (2023) utilizes a GAN to predict mel-spectrograms, while an AR model predicts phone-level prosody codes. NaturalSpeech 3 Ju et al. (2024) employs a unified framework based on discrete diffusion models to predict discrete representations of different speech attributes. However, these NAR systems need to predict phoneme-level duration, leading to a complex pipeline and more standardized generative results. SimpleSpeech Yang et al. (2024), DiTTo-TTS Lee et al. (2024), and E2 TTS Eskimez et al. (2024) are also NAR-based models that do not require precise alignment information between text and speech, nor do they predict phoneme-level duration. We discuss these concurrent works in Appendix K.
@@
Masked Generative Model. Masked generative transformers, a class of generative models, achieve significant results and demonstrate potential comparable to or superior to that of autoregressive models or diffusion models in the fields of image Chang et al. (2022, 2023); Lezama et al. (2022); Li et al. (2023a), video Yu et al. (2023a, b), and audio Borsos et al. (2023a); Garcia et al. (2023); Li et al. (2024); Ziv et al. (2024) generation. MaskGIT Chang et al. (2022) is the first work to use masked generative models for both unconditional and conditional image generation. Subsequently, Muse Chang et al. (2023) leverages rich text to achieve high-quality and diverse text-to-image generation within the same framework. MAGVIT-v2 Yu et al. (2023b) employs masked generative models with novel lookup-free quantization, outperforming diffusion models in image and video generation. Recently, some efforts have been made to adapt masked generative models to the field of audio. SoundStorm Borsos et al. (2023a) takes in the semantic tokens from AudioLM and utilizes this generative paradigm to generate tokens for a neural audio codec Zeghidour et al. (2021). VampNet Garcia et al. (2023) and MAGNeT Ziv et al. (2024) apply masked generative models for music and audio generation, while MaskSR Li et al. (2024) extends these models for speech restoration.
@@
Discrete Speech Representation. Speech representation is a crucial aspect of speech generation. Early works Ren et al. (2019); Wang et al. (2017) typically utilized mel-spectrograms as the modeling target. Recently, some large-scale TTS systems Wang et al. (2023); Ju et al. (2024) have shifted to using discrete speech representations. Discrete speech representation can be primarily divided into two types: semantic discrete representation and acoustic discrete representation [^1] We give a more detailed discussion about the definitions of “semantic” and “acoustic” in Appendix B.. Semantic discrete representations are mainly extracted from various speech SSL models Chung et al. (2021); Hsu et al. (2021); Chen et al. (2022) using quantization methods such as k-means. Acoustic discrete representations, on the other hand, are usually obtained by training a VQ-GAN model Van Den Oord et al. (2017) with the goal of waveform reconstruction, as seen in speech codecs Défossez et al. (2022); Zeghidour et al. (2021); Kumar et al. (2024). Semantic discrete representation typically shows a stronger correlation with text, whereas acoustic discrete representation more effectively reconstructs audio. Consequently, some two-stage TTS models predict both semantic and acoustic tokens. FACodec Ju et al. (2024) is a novel speech codec that disentangles speech into subspaces of different attributes, including content, prosody, timbre, and acoustic details.
@@ 3.2 Model Overview
An overview of the MaskGCT framework is presented in Figure 1. Following Betker (2023); Borsos et al. (2023a); Wang et al. (2023), MaskGCT is a two-stage TTS system. The first stage uses text to predict speech semantic representation tokens, which contain most information of content and partial information of prosody. The second stage model is trained to learn more acoustic information. Unlike previous works Betker (2023); Wang et al. (2023); Borsos et al. (2023a); Kharitonov et al. (2023) use an autoregressive model for the first stage, MaskGCT utilizes the non-autoregressive masked generative modeling paradigm for both the two stages without text-speech alignment supervision and phone-level duration prediction: (1) For the first stage model, we trained a model to learn $p_{\theta_{\text{s1}}}(\mathbf{S}|\mathbf{S}_{t},(\mathbf{S}^{p},\mathbf{P}))$, where $\mathbf{S}$ is the speech semantic representation token sequence obtained from a speech semantic representation codec (we introduce in 3.2.1), $\mathbf{S}^{p}$ is the prompt semantic token sequence, and $\mathbf{P}$ is the text token sequence. $\mathbf{S}^{p}$ and $\mathbf{P}$ are the condition for the first stage model. (2) The second stage model is trained to learn $p_{\theta_{\text{s2}}}(\mathbf{A}|\mathbf{A}_{t},(\mathbf{A}^{p},\mathbf{S}))$, where $\mathbf{A}$ is the multi-layer acoustic token sequence from a speech acoustic codec like Zeghidour et al. (2021); Défossez et al. (2022). Our second stage model is similar to SoundStorm Borsos et al. (2023a). We give more details about the four parts in the following sections.
@@
Figure 1: An overview of the proposed two-stage MaskGCT framework. It consists of four main components: (1) a speech semantic representation codec converts speech to semantic tokens; (2) a text-to-semantic model predicts semantic tokens with text and prompt semantic tokens; (3) a semantic-to-acoustic model predicts acoustic tokens conditioned on semantic tokens; (4) a speech acoustic codec reconstructs waveform from acoustic tokens.
3.2.1 Speech Semantic Representation Codec